# Regularity and stability in a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion

### v1

@@ Abstract
We consider a mathematical model of cancer cell invasion of the extracellular matrix (ECM), comprising a strongly degenerate parabolic partial differential equation for the cell volume fraction, coupled with an ordinary differential equation for the ECM volume fraction. The model captures the intricate link between the dynamics of invading cancer cells and the surrounding ECM. First, migrating cells undergo haptotaxis, i.e., movement up ECM density gradients. Secondly, cancer cells degrade the ECM fibers by means of membrane-bound proteases. Finally, their migration speed is modulated by the ECM pore sizes, resulting in the saturation or even interruption of cell motility both at high and low ECM densities. The inclusion of the physical limits of cell migration results in two regimes of degeneracy of the nonlinear diffusion and haptotaxis terms.
We devise specific estimates that are compatible with the strong degeneracy of the equation, focusing on the vacuum present at low ECM densities. Based on these regularity results, and associated local compactness, we prove stability of weak solutions with respect to perturbations on bounded initial data.
@@ 1 Introduction
Migrating cells typically move following environmental clues. Notably, they may have the ability to sense environmental gradients resulting in directed movement, for instance, down density-dependent pressure gradients or up gradients of chemical concentrations (e.g., oxygen, growth factors, etc), phenomenon known as chemotaxis. Empirical evidence also indicates the tendency of migrating cells to move by haptotaxis [3], that is up density gradients of the extracellular matrix (ECM) – i.e., the network consisting of extracellular macromolecules and elastic fibers providing cells with structural support. Haptotaxis-driven motion is particularly interesting in the context of cancer invasion, as invading cells may themselves create ECM gradients [14] by degrading the ECM fibers by means of diffusive or membrane-bound proteases [9].
Independently of the environmental clues directing cell migration, the density of the ECM plays an important role in modulating the speed at which cells can migrate through 3D tissues [23]. On one hand, invading cells rely on the presence of ECM fibers to to form focal adhesions and exert traction forces boosting them forward, and thus a lower ECM density may hinder this process. On the other hand, a denser ECM is characterised by smaller pore sizes that may slow down cell migration, by friction, or impede it altogether if the pore size is smaller than the diameter of the cell nucleus. Indeed, the proteolytic degradation of ECM fibers by cancer cells may help them achieve the optimal environmental set up for fast invasion [9].
@@
On the contrary, very few works have investigated the well-posedness of models solely considering membrane-bound proteases as in [14], which lacks the additional regularity introduced by the protease diffusivity, mainly by Zhigun and coworkers [25, 26, 24]. In these papers, the authors prove global existence of weak solutions in haptotaxis-driven cancer invasion models, with carefully defined diffusion and haptotactic coefficients to avoid blow up of solutions in finite time typical of Chaplain-Lolas inspired models [17, 19]. Interestingly, they focus on parabolic PDEs presenting a degeneracy of the diffusion in the absence of ECM density, although this degeneracy does not impact haptotaxis.
The modulation of the speed of cell migration by the ECM pore size – or, under isotropic assumptions, the ECM density or volume fraction – has been well characterised by Preziosi and coworkers [2, 10, 15]. The resulting function modelling the ECM density-dependent saturation of cell migration speed may result in the degeneracy of the overall velocity [13], both at low and high ECM density. The well-posedness of PDE models implementing such a degeneracy resulting from the physical limits of cell migration is yet to be addressed and motivates the present work.
@@
We present, in Section 2, a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion, focusing on membrane-bound proteases, that accounts for the ECM density-dependent saturation of the overall cell migration speed. In Section 3, we devise specific estimates compatible with the degeneracy of the equation. Based on the regularity provided by these estimates, we obtain some local compactness results in Section 4 and prove stability of weak solutions with respect to perturbations of bounded initial data, in Section 5. We conclude with a brief discussion in Section 6.
@@ 2 The model
${{\bm{V}}}(\phi,\varphi):=S(\varphi)\,\big{(}\alpha\nabla\varphi\,-D\nabla\psi\big{)}\,,$ (2)
@@ 3.1 Regularity results for $\varphi^{0}\leq 1$
Proposition 1 (Elementary a priori estimates)
Let $\Omega\equiv\mathbb{R}^{d}$, and let $\psi\geq 0$ and $\varphi\geq 0$ be solutions of (1) and definitions (2)-(4), complemented with initial conditions $\psi(0,{\bm{x}})=\psi^{0}({\bm{x}})$ and $\varphi(0,{\bm{x}})=\varphi^{0}({\bm{x}})$ satisfying assumptions (6). Then for all $t\in[0,\infty)$ and ${\bm{x}}\in\Omega$ we have that
@@
In addition, we have the following estimates for all $T>0$:
@@
Step 1. Upper bound on $\varphi$. The bound (8) on $\varphi$ comes directly from integrating Eq. (1)2 and using the non-negativity of $\psi$, yielding that $\varphi$ can only decrease in time. Therefore $\varphi\in L^{\infty}((0,\infty)\times\Omega)$, as in estimate (9).
@@
for some constant $C_{3}(t)$. From the same equality for the energy, integrating in time for $T\in[0,T]$ and using (15), we obtain
$DR_{0}\displaystyle\int_{0}^{T}\int_{\Omega}\psi(t,{\bm{x}})^{3}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{4}(T)\quad\text{and}\quad\int_{\Omega}\psi\varphi(1-\varphi)|\alpha\nabla\varphi-D\nabla\psi|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{5}(T),$
@@
i.e. $\psi\in L^{\infty}((0,T);L^{1}(\Omega))$. This concludes estimate (9).
@@ 3.2 Additional regularity results under assumptions (7)
Proposition 2 (Further estimates)
Let $\Omega\equiv\mathbb{R}^{d}$, and let $\psi$ and $\varphi$ be solutions of (1) and definitions (2)-(4), complemented with initial conditions $\psi(0,{\bm{x}})=\psi^{0}({\bm{x}})$ and $\varphi(0,{\bm{x}})=\varphi^{0}({\bm{x}})$ satisfying assumptions (6) and (7). We then have the following estimates for all $T>0$:
@@
$\psi\varphi(1-\varphi)|\nabla\psi|^{2}\in$ $\;L^{1}((0,T)\times\Omega),$ (19)
@@
$\frac{d}{dt}\int_{\Omega}\psi\ln(\psi)\,{\rm d}{\bm{x}}=\int_{\Omega}\nabla\psi\cdot\left[(1-\varphi)\varphi(\alpha\nabla\varphi-D\nabla\psi)\right]+\int_{\Omega}\psi(1+\ln(\psi))R_{0}(1-\psi-\varphi),$
@@
${\frac{\alpha}{\gamma}\int_{\Omega}\frac{|\nabla\varphi|^{2}}{2}(1-\varphi)\,{\rm d}{\bm{x}}\leq C_{7}(T)},\qquad D\int_{0}^{T}\int_{\Omega}(1-\varphi)\varphi|\nabla\psi|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{8}(T),\$
@@
$\int_{\Omega}|\nabla\sqrt{\varphi}|^{2}\,{\rm d}{\bm{x}}\leq C_{9}(T),\qquad\quad\int_{0}^{T}\int_{\Omega}\psi|\nabla\sqrt{\varphi}|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{10}(T)\,,$
@@
$\int_{0}^{T}\int_{\Omega}\psi\varphi(1-\varphi)|\nabla\psi|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{11}(T),$
@@
$\int_{0}^{T}\int_{\Omega}\left|\psi^{\frac{1}{2}}\;\left(\psi^{\frac{1}{2}}\nabla\sqrt{\varphi}\right)\right|^{\frac{3}{2}}\,{\rm d}{\bm{x}}{\rm d}t\leq\left(\int_{0}^{T}\int_{\Omega}\psi^{3}\,{\rm d}{\bm{x}}{\rm d}t\right)^{\frac{1}{4}}\left(\int_{0}^{T}\ \int_{\Omega}\varphi\left|\nabla\psi\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right)^{\frac{3}{4}}\,,$
@@ 4 Local compactness
Secondly, we use the Aubin-Lions-Simon lemma [18, Theorem 5] to gain time compactness of $\varphi^{\frac{1}{2}}\psi$. We combine the equations on $\varphi$ and $\psi$ in (1), to get
@@
with ${{\bm{V}}}\equiv{{\bm{V}}}(\psi,\varphi)$ and $R\equiv R(\psi,\varphi)$ given by (2)-(4). From this, to apply the Aubin-Lions-Simon lemma, it remains to estimate the right-hand-side of Eq. (22) as the divergence of a term bounded in $L^{1}_{\rm loc}$ and other terms bounded in $L^{1}_{\rm loc}$. To get this, we see that $\varphi^{\frac{1}{2}}\psi\,{{\bm{V}}}(\psi,\varphi)\in L^{1}((0,T)\times\Omega)$ from
$\int_{0}^{T}\int_{\Omega}\varphi^{\frac{1}{2}}\psi\,|{{\bm{V}}}(\psi,\varphi)|\,{\rm d}{\bm{x}}{\rm d}t$ $=\int_{0}^{T}\int_{\Omega}\left(\varphi^{\frac{1}{2}}\varphi^{\frac{1}{2}}(1-\varphi)^{\frac{1}{2}}\psi^{\frac{1}{2}}\right)\left(\psi^{\frac{1}{2}}\varphi^{\frac{1}{2}}(1-\varphi)^{\frac{1}{2}}|\alpha\nabla\varphi-D\nabla\psi|\right)\,{\rm d}{\bm{x}}{\rm d}t$
$\leq\left[\int_{0}^{T}\int_{\Omega}\psi\varphi^{2}(1-\varphi)\,{\rm d}{\bm{x}}{\rm d}t\int_{0}^{T}\int_{\Omega}\psi\varphi(1-\varphi)\left|\alpha\nabla\varphi-D\nabla\psi\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right]^{\frac{1}{2}}$
after using the Cauchy-Schwarz inequality. The first integral is bounded by (9) thanks to bound (8), and the second one is bounded by estimate (13). The second term on the right-hand side of Eq. (22) is bounded in $L^{1}$ since we have
$\int_{0}^{T}\int_{\Omega}\psi|\nabla$ $\varphi^{\frac{1}{2}}\cdot{{\bm{V}}}(\psi,\varphi)|\,{\rm d}{\bm{x}}{\rm d}t=\int_{0}^{T}\int_{\Omega}\big{|}\nabla\sqrt{\varphi}\cdot\left[\psi\varphi(1-\varphi)(\alpha\nabla\varphi-D\nabla\psi)\right]\big{|}\,{\rm d}{\bm{x}}{\rm d}t$
$\leq\left[\int_{0}^{T}\int_{\Omega}\varphi(1-\varphi)\psi\left|\nabla\sqrt{\varphi}\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\int_{0}^{T}\int_{\Omega}\psi\varphi(1-\varphi)\left|\alpha\nabla\varphi-D\nabla\psi\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right]^{\frac{1}{2}},$
again after using the Cauchy-Schwarz inequality. The first integral is bounded by estimate (18) thanks to bound (8) and under assumption (7). The second integral is bounded by estimate (13). Finally, the last two terms in Eq. (22) are bounded in $L^{1}$ thanks to (8) and (12).

### v2

# Existence, regularity and stability in a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion

@@ Abstract
We consider a mathematical model of cancer cell invasion of the extracellular matrix (ECM), comprising a strongly degenerate parabolic partial differential equation for the cell volume fraction, coupled with an ordinary differential equation for the ECM volume fraction ($0\leq\varphi\leq 1$). The model captures the intricate link between the dynamics of invading cancer cells and the surrounding ECM. First, migrating cells undergo haptotaxis, i.e., movement up ECM density gradients. Secondly, cancer cells degrade the ECM fibers by means of membrane-bound proteases. Finally, their migration speed is modulated by the ECM pore sizes, resulting in the saturation or even interruption of cell motility both at high and low ECM densities. The inclusion of the physical limits of cell migration results in two regimes of degeneracy (at $\varphi=0$ and $\varphi=1$) impacting simultaneously the nonlinear diffusion and haptotaxis terms.
We devise specific estimates that are compatible with the strong degeneracy of the equation, focusing on the vacuum present at low ECM densities. Based on these regularity results, and associated local compactness, we prove stability of weak solutions with respect to perturbations on bounded initial data, and global existence of weak solutions in the limit of an approximate non-degenerate problem.
@@ 1 Introduction
Migrating cells typically move following environmental clues. Notably, they may have the ability to sense environmental gradients resulting in directed movement, for instance, down density-dependent pressure gradients or up gradients of chemical concentrations (e.g., oxygen, growth factors, etc), phenomenon known as chemotaxis. Empirical evidence also indicates the tendency of migrating cells to move by haptotaxis [4], that is up density gradients of the extracellular matrix (ECM) – i.e., the network consisting of extracellular macromolecules and elastic fibers providing cells with structural support. Haptotaxis-driven motion is particularly interesting in the context of cancer invasion, as invading cells may themselves create ECM gradients [24] by degrading the ECM fibers by means of diffusive or membrane-bound proteases [14]. In particular, membrane-bound matrix metalloproteinases (MMPs) such as MT1-MMP, which are key to the activation of diffusive MMPs [1], have been recognised to play a prominent role in the metastatic process, with experimental findings pointing towards a direct essential role of MT-MMPs in the invasion process that is independent of the action of soluble MMPs [16, 25].
Independently of the environmental clues directing cell migration, the density of the ECM plays an important role in modulating the speed at which cells can migrate through 3D tissues [34]. On one hand, invading cells rely on the presence of ECM fibers to form focal adhesions and exert traction forces boosting them forward, and thus a lower ECM density may hinder this process. On the other hand, a denser ECM is characterised by smaller pore sizes that may slow down cell migration, by friction, or impede it altogether if the pore size is smaller than the diameter of the cell nucleus. Indeed, the proteolytic degradation of ECM fibers by cancer cells may help them achieve the optimal environmental set up for fast invasion [14].
@@
On the contrary, very few works have investigated the well-posedness of models solely considering membrane-bound proteases as in [24], which lacks the additional regularity introduced by the protease diffusivity [23], mainly by Zhigun and coworkers [35, 36, 37]. In these papers, the authors prove global existence of weak solutions in haptotaxis-driven cancer invasion models, with carefully defined diffusion and haptotactic coefficients to avoid blow up of solutions in finite time typical of Chaplain-Lolas inspired models [28, 30]. Interestingly, they focus on parabolic PDEs presenting a degeneracy of the diffusion in the absence of ECM density, although this degeneracy does not impact haptotaxis.
The modulation of the speed of cell migration by the ECM pore size – or, under isotropic assumptions, the ECM density or volume fraction – has been well characterised by Preziosi and coworkers [3, 15, 26]. The resulting function modelling the ECM density-dependent saturation of cell migration speed may result in the degeneracy of the overall velocity [22], both at low and high ECM density. The well-posedness of PDE models implementing such a degeneracy resulting from the physical limits of cell migration is yet to be addressed and motivates the present work. Naturally, the scenario of cancer invasion prominently mediated by membrane-bound MMPs, rather than diffusive ones, represents a more critical case with greater analytical challenges in dealing with the strong degeneracy of the PDE.
@@
We present, in Section 2, a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion, focusing on membrane-bound proteases, that accounts for the ECM density-dependent saturation of the overall cell migration speed. The degeneracy arises both at $\varphi=0$ and $\varphi=1$, being $\varphi$ the volume fraction occupied by the ECM, and affects both the diffusion and haptotactic terms. In Section 3, we devise specific estimates compatible with the degeneracy of the equation. In this section, we focus on the degeneracy at 0 under a stricter bound on the initial ECM volume fraction, that allows us to avoid the degeneracy at 1, which is introduced for technical reasons to obtain stronger gradient controls. Based on the regularity provided by these estimates, we obtain some local compactness results in Section 4 and prove stability of weak solutions with respect to perturbations of bounded initial data in Section 5. Existence of weak solutions is addressed in Section 6, where solutions to the degenerate problem are obtained in the limit of an approximate non-degenerate problem – see Figure 1 for a schematic illustration. We conclude with a brief discussion in Section 7, in which the biological rationale behind the stricter bound on the initial ECM volume fraction is also addressed.
@@ 2 The model
${{\bm{V}}}(\psi,\varphi):=S(\varphi)\,\big(\alpha\nabla\varphi\,-D\nabla\psi\big)\,$, (2)
@@
Figure 1: Schematic illustration of the modulation of the velocity according to the ECM volume fraction $\varphi$. Left: behaviour of the function $S(\varphi)$ defined in (3) marking the strong degeneracy of the velocity $\bm{V}$ at $\varphi=0$ and $\varphi=1$ (in red) along with the region excluded under assumption (7) (in gray). Right: corresponding illustration of the perturbed function $S_{\delta}(\varphi)=\delta+S(\varphi)>0$ used in the perturbed non-degenerate problem considered to obtain existence in the limit $\delta\to 0$ in Section 6.
@@
The major difficulty to analyse this system is the degeneracy when $\varphi(t,{\bm{x}})=0$. The degeneracy for $\varphi(t,{\bm{x}})=1$ can be avoided – see also gray area in Figure 1 – because $\varphi$ is decreasing in time and the assumption (7), which is fundamental in our analysis, is propagated with time.
@@ 3.1 Regularity results for $\varphi^{0}\leq 1$
Proposition 1 (Elementary a priori bounds)
Let $\Omega\equiv\mathbb{R}^{d}$, and let $(\psi,\varphi)$, with $\psi\geq 0$ and $\varphi\geq 0$, be a weak solution built in Section 6 of (1) and definitions (2)-(4), complemented with initial conditions $\psi(0,{\bm{x}})=\psi^{0}({\bm{x}})$ and $\varphi(0,{\bm{x}})=\varphi^{0}({\bm{x}})$ satisfying assumptions (6). Then for all $t\in[0,\infty)$ and ${\bm{x}}\in\Omega$ we have that
@@
In addition, that the following quantities are bounded in the following spaces for all $T>0$:
@@
The proposition follows from Section 6 but, in preparation, we give here the main calculations which hold for strong solutions.
Step 1. Upper bound on $\varphi$. The bound (8) on $\varphi$ comes directly from integrating Eq. (1)2 and using the non-negativity of $\psi$, yielding that $\varphi$ can only decrease in time. Therefore $\varphi\in L^{\infty}((0,\infty)\times\Omega)$.
@@
for some constant $C_{3}(t)$. From the same equality for the energy, integrating in time for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}\in[0,T]$ and using (16), we obtain
$DR_{0}\displaystyle\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi(t,{\bm{x}})^{3}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{4}(T)\quad\text{and}\quad\int_{\Omega}\psi\varphi(1-\varphi)|\alpha\nabla\varphi-D\nabla\psi|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{5}(T),$
@@
i.e. $\psi\in L^{\infty}((0,T);{L^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}}(\Omega))$. This concludes estimate (9).
@@ 3.2 Additional regularity results under assumptions (7)
Proposition 2 (Further bounds)
Let $\Omega\equiv\mathbb{R}^{d}$, and let $(\psi,\varphi)$, with $\psi\geq 0$ and $\varphi\geq 0$, be a weak solution built in Section 6 of (1) and definitions (2)-(4), complemented with initial conditions $\psi(0,{\bm{x}})=\psi^{0}({\bm{x}})$ and $\varphi(0,{\bm{x}})=\varphi^{0}({\bm{x}})$ satisfying assumptions (6) and (7). We then have the following bounds for all $T>0$:
@@
$\psi\varphi(1-\varphi)|\nabla\psi|^{2}\in$ $\;L^{1}((0,T)\times\Omega).$ (20)
Finally, we also have a bound for all $T>0$
@@
$\frac{d}{dt}\int_{\Omega}\psi\ln(\psi)\,{\rm d}{\bm{x}}=\int_{\Omega}\nabla\psi\cdot\left[(1-\varphi)\varphi(\alpha\nabla\varphi-D\nabla\psi)\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,{\rm d}{\bm{x}}}+\int_{\Omega}\psi(1+\ln(\psi))R_{0}(1-\psi-\varphi){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\,{\rm d}{\bm{x}}},$ (22)
@@
$\frac{\alpha}{\gamma}\int_{\Omega}\frac{|\nabla\varphi|^{2}}{2}(1-\varphi)\,{\rm d}{\bm{x}}\leq C_{7}(T)$$\qquad\\$ (23)
and
$D\int_{0}^{T}\hskip-5.0pt\int_{\Omega}(1-\varphi)\varphi|\nabla\psi|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{8}(T),$
@@
$\int_{\Omega}|\nabla\sqrt{\varphi}|^{2}\,{\rm d}{\bm{x}}\leq C_{9}(T),\qquad\quad\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi|\nabla\sqrt{\varphi}|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{10}(T)\,$,
@@
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi\varphi(1-\varphi)|\nabla\psi|^{2}\,{\rm d}{\bm{x}}{\rm d}t\leq C_{11}(T),$
@@
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\left|\psi^{\frac{1}{2}}\;\left(\psi^{\frac{1}{2}}\nabla\sqrt{\varphi}\right)\right|^{\frac{3}{2}}\,{\rm d}{\bm{x}}{\rm d}t\leq\left(\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi^{3}\,{\rm d}{\bm{x}}{\rm d}t\right)^{\frac{1}{4}}\left(\int_{0}^{T}\ \int_{\Omega}\varphi\left|\nabla\psi\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right)^{\frac{3}{4}}\,$,
@@ 4 Local compactness
Secondly, we use the Aubin-Lions-Simon lemma [29, Theorem 5] to gain time compactness of $\varphi^{\frac{1}{2}}\psi$. We combine the equations on $\varphi$ and $\psi$ in (1), to get
@@
with ${{\bm{V}}}\equiv{{\bm{V}}}(\psi,\varphi)$ and $R\equiv R(\psi,\varphi)$ given by (2)-(4). From this, to apply the Aubin-Lions-Simon lemma, it remains to estimate the right-hand-side of Eq. (25) as the divergence of a term bounded in $L^{1}_{\rm loc}$ and other terms bounded in $L^{1}_{\rm loc}$. To get this, we first prove that $\varphi^{\frac{1}{2}}\psi\,{{\bm{V}}}(\psi,\varphi)\in L^{1}((0,T)\times\Omega)$. This follows from
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\varphi^{\frac{1}{2}}\psi\,|{{\bm{V}}}(\psi,\varphi)|\,{\rm d}{\bm{x}}{\rm d}t$ $=\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\left(\varphi^{\frac{1}{2}}\varphi^{\frac{1}{2}}(1-\varphi)^{\frac{1}{2}}\psi^{\frac{1}{2}}\right)\left(\psi^{\frac{1}{2}}\varphi^{\frac{1}{2}}(1-\varphi)^{\frac{1}{2}}|\alpha\nabla\varphi-D\nabla\psi|\right)\,{\rm d}{\bm{x}}{\rm d}t$
$\leq\left[\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi\varphi^{2}(1-\varphi)\,{\rm d}{\bm{x}}{\rm d}t\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi\varphi(1-\varphi)\left|\alpha\nabla\varphi-D\nabla\psi\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right]^{\frac{1}{2}}$
after using the Cauchy-Schwarz inequality. The first integral is bounded by (9) thanks to bound (8), and the second one is bounded by estimate (13), allowing us to conclude a uniform bound in the space
$\varphi^{\frac{1}{2}}\psi\,|{{\bm{V}}}(\psi,\varphi)|\in L^{1}_{loc}((0,T)\times\Omega).$ (26)
The second term on the right-hand side of Eq. (25) is bounded in $L^{1}$ since we have
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi|\nabla$ $\varphi^{\frac{1}{2}}\cdot{{\bm{V}}}(\psi,\varphi)|\,{\rm d}{\bm{x}}{\rm d}t=\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\big|\nabla\sqrt{\varphi}\cdot\left[\psi\varphi(1-\varphi)(\alpha\nabla\varphi-D\nabla\psi)\right]\big|\,{\rm d}{\bm{x}}{\rm d}t$
$\leq\left[\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\varphi(1-\varphi)\psi\left|\nabla\sqrt{\varphi}\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi\varphi(1-\varphi)\left|\alpha\nabla\varphi-D\nabla\psi\right|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right]^{\frac{1}{2}}\hskip-3.0pt,$
again after using the Cauchy-Schwarz inequality. The first integral is bounded by estimate (19) thanks to bound (8) and under assumption (7). The second integral is bounded by estimate (13). Finally, the last two terms in Eq. (25) are bounded in $L^{1}$ thanks to (8) and (12). These bounds conclude time compactness proof.